6(X-2)+2X=8(X+2)-28
X=?

Found 2 solutions by Boreal, MathTherapy:
Answer by Boreal(15235)

6(X-2)+2X=8(X+2)-28
6x-12 +2x=8x-16-28
4x-12 =8x -44
-4x =-32
x=8 ANSWER
check
6(6)+2(8)=52
8(10)-28 =52

Answer by MathTherapy(10551)

6(X-2)+2X=8(X+2)-28
X=?
Left side of equation = Right side of equation.
Therefore, any value for x will be a solution.